In a constant volume gas thermometer the gas pressure at 0 degree Celsius is 200mmHg and at 100 degree Celsius it is 270mmHg what is the temperature if the thermometer produces a pressure of 214mmHg


1
Expert's answer
2021-12-27T13:02:29-0500

Solution:


"p_1=200mmHg;"

"t_1=0^oC;"

"p_2=270mmHg;"

"t_2=100^oC;"

"V=const;"

"p=214mmHg;"

"t-?;"

what is the temperature if the thermometer produces a pressure of 214mmHg:

"\dfrac{\Delta p}{\Delta t}=\dfrac{p_2-p_1}{t_2-t_1};"

"\dfrac{\Delta p}{\Delta t}=\dfrac{270mmHg-200mmHg}{100^oC-0^oC}=0.7\dfrac{mmHg}{^oC};"

"\dfrac{\Delta p}{\Delta t}=\dfrac{p-p_1}{t-t_1};\implies t-t_1=\dfrac{p-p_1}{\dfrac{\Delta p}{\Delta t}};"

"t-t_1=\dfrac{214mmHg-200mmHg}{0.7\dfrac{mmHg}{^oC}}=20^oC;"

"t=20^oC+0^oC=20^oC;"

Answer:


 "t=20^oC;"